Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $127,400 of manufacturing overhead for an estimated allocation base of $91,000 direct material dollars to be used in production. The company has provided the following data for the just completed year:

Purchase of raw materials $ 139,000
Direct labor cost $ 87,000
Manufacturing overhead costs:
Indirect labor $ 127,900
Property taxes $ 9,000
Depreciation of equipment $ 17,000
Maintenance $ 15,000
Insurance $ 9,400
Rent, building $ 35,000
Beginning Ending
Raw Materials $ 23,000 $ 13,000
Work in Process $ 44,000 $ 35,000
Finished Goods $ 73,000 $ 59,000

Required:

1. Compute the predetermined overhead rate for the year.

2. Compute the amount of underapplied or overapplied overhead for the year.

3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

5. Assume that the $35,000 ending balance in Work in Process includes $8,800 of direct materials. Given this assumption, supply the information missing below:
Answer #1

1) Predetermine overhead rate = 127400/91000 = 140% of direct material used

2) Applied overhead = (23000+139000-13000)*1.4 = 208600

Actual overhead = 213300

Under applied overhead = 213300-208600 = 4700

3) Schedule of cost of goods manufactured

Direct material
Beginning raw material 23000
Raw material purchase 139000
Raw material available for use 162000
Less: Ending raw material -13000
Direct material used 149000
Direct labor 87000
Manufacturing overhead 208600
Total manufacturing cost 444600
Beginning work in process 44000
Total cost of work in process 488600
Less: Ending work in process -35000
Cost of goods manufactured 453600

4) Unadjusted cost of goods sold = 73000+453600-59000 = 467600

5) Calculate following

Direct material 8800
Direct labor 13880
Overhead (8800*1.4) 12320
Ending WIP 35000